# Image Matching Filtering and Refinement by Planes and Beyond

### v3

@@ 3.2.4 Fixing rotations
MiHo is not rotation invariant. Figure 3 illustrates how midpoints change with rotations. In the ideal MiHo configuration where the images are upright the area of the image formed in the midpoint plane is within the areas of the original ones, making a single cone when considering the images as stacked. The area of the midpoint plane is lower than the minimum area between the two images when there is a rotation about , and could degenerate to the apex of the corresponding double cone. A simple strategy to get an almost-optimal configuration was experimentally verified to work well in practice under the observation that images acquired by cameras tend to have a possible relative rotation of a multiple of .
where
and is the Iverson bracket. The global orientation estimation can be computed efficiently on the initial input matches before running MOP, adjusting keypoints accordingly. Final homographies can be adjusted in turn by removing the rotation given by . Figure 4 shows the MOP+MiHo result obtained on a set of matches without or with orientation pre-processing, highlighting the better matches obtained in the latter case.
@@ 3.3 Normalized Cross Correlation (NCC)
3.3.1 Local image warping
NCC is employed to effectively refine corresponding keypoint localization. The approach assumes that corresponding keypoints have been roughly aligned locally by the transformations provided as a homography pair. In the warped aligning space, template matching [48] is employed to refine translation offsets in order to maximize the NCC peak response when the patch of a keypoint is employed as a filter onto the other image of the corresponding keypoint.
@@ 3.3.1 Local image warping
and the other is extended pair defined as
@@
as described in Sec. 3.2.1, in case of MOP+MiHo
@@
3.3.2 Computation
such that the squared window set of offsets is
and has cardinality equal to
The squared patch of centered in with radius is the pixel set
while the mean intensity value of the patch is
and the variance is
which is ideally robust to affine illumination changes [48]. Lastly, the similarity between two patches and is
so that, for the match and a set of associated homography pairs derived in Eq. 3.3.1, the refined keypoint offsets and the best aligning homography pair are given by
is obtained by reprojecting back to the original images, i.e.
3.3.3 Sub-pixel precision
The refinement offset can be further enhanced by parabolic interpolation adding sub-pixel precision [102]. Notice that other forms of interpolations have been investigated but they did not provide any sub-pixel accuracy improvements according to previous experiments [49]. The NCC response map at in the best warped space with origin the peak value can be written as
from computed by Eq. 3.3.2. The sub-pixel refinement offsets in the horizontal direction are computed as the vertex of a parabola fitted on the horizontal neighborhood centered in the peak of the maximized NNC response map, i.e.
and from Eq. 3.3.2 the final refined match
becomes
@@ 4.1.1 Base pipelines
Seven base pipelines have been tested to provide the input matches to be filtered. For each of these, the maximum number of keypoints was set to 8000 to retain as many as possible matches. This leads to including more outliers and better assessing the filtering potential. Moreover, nowadays benchmarks and deep matching architecture designs assume upright image pairs, where the relative orientation between images is roughly no more than . This additional constraint allows us to retrieve better initial matches in common user acquisition scenarios. Current general standard benchmark datasets, including those employed in this evaluation, are built so that the image pairs do not violate this constraint. Notice that many SOTA joint and end-to-end deep architectures do not tolerate strong rotations within images by design.
SIFT+NNR [19] is included as the standard and reference handcrafted pipeline. The OpenCV [103] implementation was used for SIFT, exploiting RootSIFT [104] for descriptor normalization, while NNR implementation is provided by Kornia [105]. To stress the match filtering robustness of the successive steps of the matching pipeline, which is the goal of this evaluation, the NNR ratio threshold was set rather high, i.e. to 0.95, while common adopted values range in depending on the scene complexity, with higher values yielding to less discriminative matches and then possible outliers. Moreover, upright patches are employed by zeroing the dominant orientation for a fair comparison with other recent deep approaches.
Key.Net++NNR [11], a modular pipeline that achieves good results in common benchmarks, is also taken into account. Excluding the NNR stage, it can be considered a modular deep pipeline. The Kornia implementation is used for the evaluation. As for SIFT, the NNR threshold is set very high, i.e. to 0.99, while more common values range in . The deep orientation estimation module OriNet [84] generally siding AffNet is not included to provide upright matches.
Other base pipelines considered are SOTA full end-to-end matching networks. In details, these are SuperPoint+LightGlue, DISK+LightGlue and ALIKED+LightGlue as implemented in [12]. The input images are rescaled so the smaller size is 1024 following LightGlue default. For ALIKED the aliked-n16rot model weights are employed which according to the authors [13] can handle slight image rotation better and are more stable on viewpoint change.
The last base pipelines added in the evaluation are DeDoDe v2 [93], which provides an alternative end-to-end deep architecture different from the above, and LoFTR, a semi-dense detector-free end-to-end network. These latter pipelines also achieve SOTA results in popular benchmarks. The authors’ implementation is employed for DeDoDe v2, setting to matching threshold to 0.01, while LoFTR implementation available through Kornia is used in the latter case.
@@ 4.1.2 Match filtering and post-processing
Eleven match filters are applied after the base pipelines and included in the evaluation besides MOP+MiHo+NCC. For the proposed filtering sub-pipeline, all the five available combinations have been tested, i.e. MOP, MOP+NCC, NCC, MOP+MiHo, MOP+MiHo+NCC, to analyze the behavior of each module. In particular, applying NCC after SIFT+NNR or Key.Net++NNR can highlight the improvement introduced by MOP which uses homography-based patch normalization against similarity or affine patches, respectively. Likewise, applying NCC to the remaining end-to-end architectures should remark the importance of patch normalization. The setup parameters indicated in Sec. 3 is employed, which has worked well in preliminary experiments. Notice that MiHo orientation adjustment described in Sec. 3.2.4 is applied although not required in the case of upright images. This allows us to indirectly assess the correctness of base assumptions and the overall general method robustness since in the case of wrong orientation estimation the consequential failure can only decrease the match quality.
To provide a fair and general comparison, the analysis was restricted to filters that require as input only the keypoint positions of the matches and the image pair. This excludes approaches requiring in addition descriptor similarities [25, 28, 27], related patch transformations [71], intermediate network representations [42, 43] or a SfM framework [101]. The compared methods are GMS [26], AdaLAM [33] and CC [79] as handcrafted filters, and ACNe [35], CLNet [37], ConvMatch [40], DeMatch [41], FC-GNN [44], MS2DG-Net [39], NCMNet [38] and OANet [36] as deep modules. The implementations from the respective authors have been employed for all filters with the default setup parameters if not stated otherwise. Except FC-GNN and OANet, deep filters require as input the intrinsic camera matrices of the two images, which is not commonly available in real practical scenarios the proposed evaluation was designed for. To bypass this restriction, the approach presented in ACNe was employed, for which the intrinsic camera matrix
@@
The above focal length estimation is quite reasonable by the statistics reported in Fig. 5 that have been extracted from the evaluation data and discussed in Sec. 4.1.3. Notice also that most of the deep filters have been trained on SIFT matches, yet to be robust and generalizable they must be able to mainly depend only on the scene and not on the kind of feature extracted by the pipeline.
For AdaLAM, no initial seeds or local similarity or affine clues were used as additional input as discussed above. Nevertheless, this kind of data could have been used only for SIFT+NNR or Key.Net++NNR. For CC, the inliers threshold was experimental set to px as in Sec. 3.1.2 instead of the value 1.5 px. suggested by authors. Otherwise, CC would not be able to work in general scenes with no planes. Note that AdaLAM and CC are the closest approaches to the proposed MOP+MiHo filtering pipeline. Moreover, among the compared methods, only FC-GNN is able to refine matches as MOP+MiHo+NCC does.
RANSAC is also optionally evaluated as the final post-process to filtering matches according to epipolar or planar constraints. Three different RANSAC implementations were initially considered, i.e. the RANSAC implementation in PoseLib333https://github.com/PoseLib/PoseLib, DegenSAC444https://github.com/ducha-aiki/pydegensac [24] and MAGSAC555https://github.com/danini/magsac [23]. The maximum number of iterations for all RANSACs is set to , which is uncommonly high, to provide a more robust pose estimation and to make the whole pipeline only depend on the previous filtering and refinement stages. Five different inlier threshold values , i.e.
are considered. After a preliminary RANSAC ablation study666https://github.com/fb82/MiHo/tree/main/data/results according to proposed benchmark, the best general choice found is MAGSAC with 0.75 px. or 1 px. as inlier threshold , indicated by MAGSAC↓ and MAGSAC↑ respectively. On one hand, the former and more strict threshold is generally slightly better in terms of accuracy. On the other hand, the latter provides more inliers, so both results will be presented. Matching outputs without RANSAC are also reported for a complete analysis.
@@ 4.1.3 Datasets
Four different datasets have been employed in the evaluation. These include the MegaDepth [54] and ScanNet [55] datasets, respectively outdoor and indoor. The same 1500 image pairs for each dataset and Ground-Truth (GT) data indicated in the protocol employed in [14] are used. These datasets are de-facto standard in nowadays image matching benchmarking; sample image pairs for each dataset are shown in Fig. 1. MegaDepth test image pairs belong to only 2 different scenes and are resized proportionally so that the maximum size is equal to 1200 px, while ScanNet image pairs belong to 100 different scenes and are resized to px. Although according to previous evaluation [12] LightGlue provides better results when the images are rescaled so that the maximum dimension is 1600 px, in this evaluation the original 1200 px resize was maintained since more outliers and minor keypoint precision are achieved in the latter case, providing a configuration in which the filtering and refinement of the matches can be better revealed. For deep methods, the best weights that fit the kind of scene, i.e. indoor or outdoor, are used when available.
To provide further insight into the case of outdoor scenes the Image Matching Challenge (IMC) PhotoTourism dataset777https://www.kaggle.com/competitions/image-matching-challenge-2022/data is also employed. The IMC PhotoTourism (IMC-PT) dataset is a curated collection of sixteen scenes derived by the SfM models of the Yahoo Flickr Creative Commons 100 Million (YFCC100M) dataset [106]. For each scene 800 image pairs have been randomly chosen, resulting in a total of 12800 image pairs. These scenes also provide 3D model scale as GT data so that metric evaluations are possible. Note that MegaDepth data are is roughly a subset of IMC-PT. Furthermore, YFCC100M has been often exploited to train deep image matching methods so it can be assumed that some of the compared deep filters are advantaged on this dataset, but this information cannot be retrieved or elaborated. Nevertheless, the proposed modules are handcrafted and not positively biased so the comparison is still valid for the main aim of the evaluation.
Lastly, the Planar dataset contains 135 image pairs from 35 scenes collected from HPatches [56], the Extreme View dataset (EVD) [57] and further datasets aggregated from [70, 32, 50]. Each scene usually includes five image pairs except those belonging to EVD consisting of a single image pair. The planar dataset includes scenes with challenging viewpoint changes possibly paired with strong illumination variations. All image pairs are adjusted to be roughly non-upright. Outdoor model weights are preferred for deep modules in the case of planar scenes. A thumbnail gallery for the scenes of the Planar dataset is shown in Fig. 6.
@@ 4.1.4 Error metrics
For non-planar datasets, unlike most common benchmarks [14], it is assumed that no GT camera intrinsic parameters are available. This is the general setting that occurs in most practical SfM scenarios, especially in outdoor environments. For a given image pair, the fundamental matrix can be estimated as
where are respectively the essential matrix, the rotation matrix, the intrinsic camera matrices and the skew-symmetric matrix corresponding to the translation vector [45]. The maximum epipolar error for a match according to is
@@
Given a set of matches obtained by a base pipeline of those presented in Sec. 4.1.1 and the final refined match set according to one of the successive post-processing listed in Sec. 4.1.2, the GT fundamental matrix is estimated according to Eq. 51 from the GT intrinsic and extrinsic camera parameters888actually these are quasi or pseudo GT data estimated from wide and reliable SfM models. in the case of the non-planar scene, or the GT homography is provided for planar scenes. The notation indicates the usage of the GT matrix according to the scene kind. The recall is computed as
@@
limited by . As additional statistic measurement the number of filtered matches is computed as
For a zero denominator in Eq. 53 the recall is forced to 0. Notice that in the case of a nonzero denominator, the recall is less than or equal to 1 only when which does not hold for pipelines where the keypoint refinement step effectively improves the matches. The number of filtered matches and the precision are less than or equal to 1 in any case by pipeline design. The final aggregated measurement for each error and dataset is provided by averaging on all dataset image pairs. It should be remarked that for non-planar scenes both precision and recall are non-perfect measures since epipolar error constraints can hold also for outliers in specific and common scene configurations [32].
To provide more reliable measures also in view of the final aim of the image matching, the relative pose error is employed for the non-planar scenes. Two alternative estimations of the pose are given in terms of rotation and translation matrices. In the first one, the fundamental matrix is estimated from the matches in by the 8-point algorithm using the normalized Direct Linear Transformation (DLT) [45] by the OpenCV function findFundamentalMat as and then the essential matrix is obtained as
given the GT intrinsic cameras and . The four possible choice of camera rotation and translation pairs are then obtained by SVD decomposition [45] using the OpenCV decomposeEssentialMat function as
@@
the essential matrix is computed directly as by the OpenCV function findEssentialMat using the 5-point algorithm [107] embedded into RANSAC with a normalized threshold [14] and the maximim number of iteration equal to . The two possible solutions for the camera rotation and translation components are recovered by the OpenCV function recoverPose as
Notice that this second approach makes it easier to find a correct solution since five points are required instead of eight for the minimum model and and are used earlier stage of the process. Nevertheless, these conditions are more unlikely to be met in practical real situations, especially in outdoor environments. Notice that the translation vectors in both and are defined up to scale. The final pose accuracy is provided in terms of the Area Under the Curve (AUC) of the best maximum angular error estimation for each dataset image pair and a given angular threshold
@@
with indicating the use of one of the two options between Eq. 58 and Eq. 60 for the estimated pose components, and being the respective GT translation, defined again up to scale, and rotation. In the above formula, the angular error for the translation and rotation are respectively computed as
@@
The translation thresholds are reasonable since the 3D scene extensions are roughly of the order of 100 m. This metric pose accuracy estimation provides a reasonable choice able to discriminate when the real baseline is negligible with respect to the whole scene so that the viewpoint change appears as rotation only. In this case, the translation angular component can be quite high as noise but it does not affect the pose estimation effectively. Using Eq. 63 would stress this error without a real justification [108]. Likewise above, the corresponding AUC values for the chosen pose estimation between Eq. 58 and Eq. 60 will be denoted by AUC, AUC respectively, for the given threshold pair . The final metric accuracy scores are given by the average over the value pairs and indicated AUC and AUC, respectively, where the ‘’ subscript indicates the metric translation error according to Eq. 65.
where the average reprojection error in a single image is estimated as
The AUC of the maximum average reprojection error in the common area for each image pair in the dataset and a fixed angular threshold
is denoted by AUC. The final homography accuracy is provided as the mean AUC over the different thresholds and is indicated by AUC to be consistent with the previous AUC notation adopted in the case of non-planar images, where the superscript ‘’ clearly stands for homography. Accuracy estimation considering the common area is more reliable than an accuracy estimation by averaging the reprojection error only in the four image corners as in [14]. This can be understood by observing that the GT planar homography is a pseudo GT estimated from correspondences, which lie inside the boundaries of both images. As an image point gets reprojected in the other image outside and far from the boundaries, which generally happens with the image corners as depicted in Fig. 2, there is no way to check the real error, which also increases going away from the keypoint matches being employed to estimate the GT homography.

### v4

@@ 3.2.5 Fixing rotations
MiHo is not rotation invariant. Figure 4 illustrates how midpoints change with rotations. In the ideal MiHo configuration, where the images are upright, the area of the image formed in the midpoint plane is within the areas of the original ones, making a single cone when considering the images as stacked. The area of the midpoint plane is lower than the minimum area between the two images when there is a rotation about , and could degenerate to the apex of the corresponding double cone. In analogous way, reflections in the worst case degenerate into segments.
Notice that a similar behavior can be observed in VSAC half homography. More in detail, the matrix entries of the square root decomposition of the original homography must not be complex. This implies that must be in case multiplied by -1 if its determinant is negative since otherwise . After this adjustment, all eigenvalues must be positive, so that reflections or rotation by cannot be handled, as well as any transformation which can be represented by a triangular matrix with one or two negative diagonal elements, including hence the related affine transformations.
A simple strategy to get an almost-optimal configuration with MiHo was experimentally verified to work well in practice under the observation that images acquired by cameras tend to have relative rotations which are multiples of . The main idea is to choose the global rotation which maximizes the configuration where midpoint inter-distances are between corresponding keypoint inter-distances and to adjust the images accordingly. This can be computed efficiently and can be extended to take into account also reflections. An analogous approach can be applied to VSAC. Further details can be found in Appendix A.
Finally, it is also possible to directly excluding reflections in the RANSAC loop inside MOP. This further variation is denoted by MOP⋆ and detailed in Appendix B.
@@ 3.3 Normalized Cross Correlation (NCC)
NCC is employed to effectively refine corresponding keypoint localization. The approach assumes that corresponding keypoint patches have been roughly aligned locally by the transformations provided as a homography pair, for instance, by MOP+MiHo. Indeed, each homography pair provides a consistent canonical frame for the patches associated to a match. In the warped aligning space, template matching [gonzales_wood] is employed to refine translation offsets in order to maximize the NCC peak response when the patch of a keypoint is employed as a filter onto the other image of the corresponding keypoint.
@@
and the other is the extended pair defined as
@@
as described in Sec. 3.2.1 in case of MOP+MiHo;
as described in Sec. 3.2.3 in case of MOP+VSAC;
@@
For each homography pair in , the associated patches are warped accordingly and NCC is used to get the pixel offset which identify the estimate discrete refinement. The keypoint offset corresponding to the homography pair globally maximizing the NCC similarity is finally selected and refined with sub-pixel precision by parabolic interpolation. Please refer to Appendix C for more details.
It is well known that NCC template matches is not suitable in the case of flat, homogeneous patches or in the presence of edges [zeliski_book]. NCC can be extended to alleviate this issue, by limiting the offset search through a mask obtained by inscribing a maximum ellipsis with the shape provided by the gradient autocorrelation matrix on the squared patch that defines NCC search space. In this way, for instance, if the patch content is just an edge in a particular direction, NCC refinements are mainly attempted along the orthogonal edge direction, where there is an effective confidence of performing this operation. This extended NCC will be denoted by NCC⋆.
@@ 4.1.1 Base pipelines
Seven base pipelines have been tested to provide the input matches to be filtered. These include SIFT+NNR [sift] as the standard handcrafted reference, Key.Net++NNR [imw2020] as a modular almost deep pipeline, SuperPoint+LightGlue, DISK+LightGlue and ALIKED+LightGlue [lightglue] as sparse fully end-to-end deep image matching architectures, DeDoDe v2 [dedodev2] which provides an alternative end-to-end deep network different from the previous ones, and LoFTR, a semi-dense detector-free end-to-end network. Additionally, further specific experiments have been carried out with the recent MASt3R [mast3r] and RoMa [roma] dense image matching architectures, for a total of nine base pipelines.
The maximum number of keypoints was set to 8000 to retain as many as possible matches, with the exception of more recent MASt3R and RoMa. This leads to including more outliers and better assessing the filtering potential. Moreover, nowadays benchmarks and deep matching architecture designs assume upright image pairs, where the relative orientation between images is roughly no more than . This additional constraint allows us to retrieve better initial matches in common user acquisition scenarios. Current general standard benchmark datasets, including those employed in this evaluation, are built so that the image pairs do not violate this constraint. Notice that many SOTA joint and end-to-end deep architectures do not tolerate strong rotations within images by design. Finally, for both MASt3R and RoMa the number of maximum output matches was set to 2048, considering computation resource constraints. Further setup details for each specific pipeline can be found in Appendix D.1.
@@ 4.1.2 Match filtering and post-processing
Besides MOP+MiHo+NCC, Thirteen match filters are applied after the base pipelines and included in the evaluation. For the proposed filtering sub-pipeline, all the five available combinations have been tested, i.e. MOP, MOP+NCC, NCC, MOP+MiHo, MOP+MiHo+NCC, to analyze the behavior of each module. In particular, applying NCC after SIFT+NNR or Key.Net++NNR can highlight the improvement introduced by MOP which uses homography-based patch normalization against similarity or affine patches, respectively. Likewise, applying NCC to the remaining end-to-end architectures should remark the importance of patch normalization. The setup parameters indicated in Sec. 3 are employed, which have worked well in preliminary experiments. Notice that MiHo orientation adjustment described in Sec. 3.2.5 is applied although not required in the case of upright images. This would indirectly assess the correctness of base assumptions and the overall general method robustness because in the case of wrong orientation estimation the consequential failure can only decrease the match quality.
In order to further analyze the behavior and robustness of the proposed approaches, additional ablation studies replacing MiHo with VSAC, as discussed in Sec. 3.2.3, and NCC with NCC⋆, as detailed in Sec. 3.3, are also included in the evaluation. In this case the MOP⋆ implementation defined in Sec. 3.2.5 has been used to prevent issues with planar image reflections in RANSAC.
To provide a fair and general comparison, the analysis was restricted to methods that require as input only the keypoint positions of the matches and the image pair. This excludes approaches requiring in addition descriptor similarities [lpm, pfm, pgm], related patch transformations [sift_affine], specific non-general pipelines [patch2pix] or a SfM framework [pixelperfectsfm]. The compared methods are GMS [gms], AdaLAM [adalam], CC [cc] and Progressive-X [progx] as handcrafted filters, and ACNe [acne], CLNet [clnet], ConvMatch [convmatch], DeMatch [dematch], FC-GNN [fcgnn], MS2DG-Net [ms2dgnet], NCMNet [ncmnet], OANet [oanet] and Keypt2Subpx [keypt2subpx] as deep modules. The implementations from the respective authors have been employed for all methods with the default setup parameters if not stated otherwise. Keypt2Subpx is only evaluated with SuperPoint+LightGlue since there are no pretrained models available for the other specific base pipeline employed in this evaluation.
With the exception of FC-GNN, Keypt2Subpx and OANet, deep filters require as input the intrinsic camera matrices of the two images, which is not commonly available in real practical scenarios the proposed evaluation was designed for. To bypass this restriction, the approach presented in ACNe was employed, for which the intrinsic camera matrix
@@
The above focal length estimation is quite reasonable according to the statistics reported in Appendix D.2 that have been extracted from the evaluation data, discussed in Sec. 4.1.3. Notice also that most of the deep filters have been trained on SIFT matches, yet to be robust and generalizable they must be able to depend mainly only on the scene and not on the kind of feature extracted by the pipeline.
For AdaLAM, no initial seeds or local similarity or affine clues were used as additional input as discussed above. Nevertheless, this kind of data could have been used only for SIFT+NNR or Key.Net++NNR. For CC and Progressive-X, the inliers threshold was experimental set to px, as defined for MOP in Sec. 3.1.2, instead of the value of 1.5 px suggested by authors. Otherwise, both filters would not be able to work in general scenes with no planes. Note that AdaLAM, CC and Progressive-X are the closest approaches to the proposed MOP+MiHo filtering pipeline. Moreover, among the compared methods, only FC-GNN and Keypt2Subpx are able to refine matches, and only FC-GNN is able to both filtering and refining matches as MOP+MiHo+NCC does.
RANSAC is also optionally evaluated as the final post-process step to filter matches according to epipolar or planar constraints. Three different RANSAC implementations were initially considered, i.e. the RANSAC implementation in PoseLib333https://github.com/PoseLib/PoseLib, DegenSAC444https://github.com/ducha-aiki/pydegensac [degensac] and MAGSAC555https://github.com/danini/magsac [magsac]. The maximum number of iterations for all RANSACs was set to , which is uncommonly high, in order to provide a more robust pose estimation and to make the whole pipeline mostly depending on the previous filtering and refinement stages. Five different inlier threshold values , i.e.
were considered. After a preliminary RANSAC ablation study666https://github.com/fb82/MiHo/tree/main/data/results according to proposed benchmark, the best general choice found is MAGSAC with px, followed by MAGSAC with px. On one hand, the former and more strict threshold is generally slightly better in terms of accuracy. On the other hand, the latter provides more inliers. MAGSAC results with px are provided as additional material for most of the evaluation configurations. Matching outputs without RANSAC are also reported to complete the analysis.
@@ 4.1.3 Datasets
Four different datasets have been employed in the evaluation, detailed descriptions can be found in Appendix D.2. Non-planar datasets include MegaDepth [megadepth] and ScanNet [scannet], respectively outdoor and indoor, together with the IMC PhotoTourism (IMC-PT) dataset777https://www.kaggle.com/competitions/image-matching-challenge-2022/data which provides further insight in the case of outdoor scenes. Lastly, the Planar dataset contains around 140 image pairs collected from HPatches [hpatches], the Extreme View dataset (EVD) [mods] and further datasets aggregated from [hess_lapl_affine, dtm, slime].
@@ 4.1.4 Error metrics
For non-planar datasets, unlike most common benchmarks [loftr], it is assumed that no Ground-Truth (GT) camera intrinsic parameters are available. This is the general setting that occurs in most practical SfM scenarios, especially in outdoor environments. For a given image pair, the fundamental matrix can be estimated as
where are respectively the essential matrix, the rotation matrix, the intrinsic camera matrices and the skew-symmetric matrix corresponding to the translation vector [multiview]. The maximum epipolar error for a match according to is
@@
Given a set of matches , obtained by a base pipeline of those presented in Sec. 4.1.1, and the final refined match set according to one of the successive post-processing listed in Sec. 4.1.2, the GT fundamental matrix is estimated by Eq. 32 from the GT intrinsic and extrinsic camera parameters888actually these are quasi or pseudo GT data estimated from wide and reliable SfM models. in the case of the non-planar scene, or the GT homography is provided for planar scenes. The notation indicates the usage of the GT matrix according to the scene kind. The recall is computed as
@@
limited by . As additional statistics, the number of filtered matches is computed as
For a zero denominator in Eq. 34 the recall is forced to 0. Notice that in the case of a non-zero denominator, the recall is less than or equal to 1 only when which does not hold for pipelines where the keypoint refinement step effectively improves the matches. The number of filtered matches and the precision are less than or equal to 1 in any case by pipeline design. The final aggregated measurement for each error and dataset is provided by averaging on all dataset image pairs. It should be remarked that for non-planar scenes both precision and recall are non-perfect measures since epipolar error constraints can hold also for outliers in specific and common scene configurations [dtm].
To provide more reliable measures also in view of the final aim of the image matching, the relative pose error is employed for the non-planar scenes. Two alternative estimations of the pose are given in terms of rotation and translation matrices. In the first one, the fundamental matrix is estimated from the matches in by the 8-point algorithm using the normalized Direct Linear Transformation (DLT) [multiview] by the OpenCV function findFundamentalMat as and then the essential matrix is obtained as
given the GT intrinsic cameras and . The four possible choice of camera rotation and translation pairs are then obtained by SVD decomposition [multiview] using the OpenCV decomposeEssentialMat function as
@@
the essential matrix is computed directly as by the OpenCV function findEssentialMat using the 5-point algorithm [nister] embedded into RANSAC with a normalized threshold [loftr] and the maximum number of iteration equal to . The two possible solutions for the camera rotation and translation components are recovered by the OpenCV function recoverPose as
Notice that this second approach makes it easier to find a correct solution since five points are required instead of eight for the minimum model and and are used earlier stage of the process. Nevertheless, these conditions are more unlikely to be met in practical real situations, especially in outdoor environments. Note that the translation vectors in both and are defined up to scale. The final pose accuracy is provided in terms of the Area Under the Curve (AUC) of the best maximum angular error estimation for each dataset image pair and a given angular threshold
@@
with indicating the use of one of the two options between Eq. 39 and Eq. 41, and being the respective GT translation, defined again up to scale, and rotation. In the above formula, the angular error for the translation and rotation are respectively computed as
@@
In the case of IMC-PT dataset, a fully metric estimation of the translation component of the pose is possible since the real GT reconstruction scale is available, i.e. is the real metric measurement. Following the IMC 2022 benchmark7, the angular thresholds of Eq. 42 are replaced by pairs of thresholds , where is a translation thresholds. The corresponding AUC values for the chosen pose estimation between Eq. 39 and Eq. 41 will be denoted by AUC, AUC respectively, for the given threshold pair . The final metric accuracy scores are given by the average over the value pairs and denoted as AUC and AUC, respectively, where the ‘’ subscript indicates the metric translation error.
For the planar dataset, the homography accuracy is measured analogously by AUC, replacing the pose error by the maximum of the average reprojection error in the common area when using one of the two images as a reference, leading to the AUC value for a given threshold . The final homography accuracy is provided as the mean AUC over the different thresholds and is indicated by AUC to be consistent with the previous AUC notation adopted in the case of non-planar images, where the superscript ‘’ clearly stands for homography.
Detailed descriptions on both the fully metric and planar AUC estimations can be found in Appendix D.3.